In the cushion of a couch, you find 22 coins, all nickels and dimes, totaling $1.70. How many of each coin do you have?
Answer by Boreal(15235)

x+y=22, where x=nickels and y=dimes
.05x+.10y=1.70; multiply this by 100 to clear decimals
5x+10y=170
-5x-5y=-110, multiplying the top by -5
5y=60
y=12 dimes or $1.20
x=10 nickels or $0.50
They add to $1.70